What is the kinetic energy of a 1.5x10^3 kg (1500 kg using sig figs) car traveling at 25 m/s?


1
Expert's answer
2021-02-28T07:36:04-0500
"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot1500\ kg\cdot(25\ \dfrac{m}{s})^2=4.68\cdot10^5\ J."